A vehicle accelerates from rest to 24.6m/s at a constant 1.98 m/s^2. What distance did the vehicle cover during this acceleration?
1
Expert's answer
2020-10-13T09:36:39-0400

We can find the distance covered by the vehicle from the kinematic equation:


"v_f^2 = v_i^2 + 2as,""s = \dfrac{v_f^2-v_i^2}{2a}=\dfrac{(24.6 \ \dfrac{m}{s})^2-(0 \ \dfrac{m}{s})^2}{2 \cdot 1.98 \ \dfrac{m}{s^2}} = 153 \ m."

Answer:

"s = 153 \ m."